### v1

@@ 4.2 Exponential error trust update
[…]
@@ 4.3 Variable-share error trust update
The variable-share error trust update method extends the exponential error trust update by introducing a redistribution mechanism that dynamically adjusts trust values based on observed prediction errors and redistributes a fraction of trust among the sources (Herbster and Warmuth 1998). This approach balances penalizing poorly performing sources and rewarding better performing ones by incorporating a flexible share allocation component.
@@
$W^{(i)}_{\text{trust}}:=\sum_{h\in[H]}(1-(1-\beta)^{\bm{v}^{(i)}_{h}})\cdot(t^{\prime})^{(i)}_{h},$
@@ 4.4 Analyzing trust dynamics
In this section, we are interested in analyzing the dynamics of trust assigned to different sources under specific trust update algorithms and understanding whether and how dominant trust exists. We focus on two methods: the min-max error trust update in Section 4.1 and the exponential error trust update in Section 4.2. We exclude the variable-share error update method in Section 4.3 from the analysis because its primary objective is to rapidly adjust trust levels when relative accuracy relationships between sources change. Consequently, the concept of dominant trust is less relevant to the variable-share error update method.
@@ 4.4.1 Dominant trust under min-max error trust update
We consider the concept of probability dominance (Wrather and Yu 1982) to rationalize the assignment of dominant trust (if it exists) to one source over others. Let $\Omega$ denote a collection of random variables corresponding to the deviations of the predicted values from the true realizations, specifically the norm of the prediction errors. For simplicity, we first examine the case where there are only two data sources, $h_{Y}$ and $h_{Z}$, that is, $H=2$. Let $Y=\|\Delta\bm{\xi}^{(i)}_{h_{Y}}\|$ represent the norm of the prediction errors of source $h_{Y}$, and $Z=\|\Delta\bm{\xi}^{(i)}_{h_{Z}}\|$ for the source $h_{Z}$, over multiple events $i=1,2,\ldots$. The cumulative distribution functions of $Y$ and $Z$ are well defined and are denoted by $F_{Y}(a)$ and $F_{Z}(a)$, respectively. In this framework, a source that consistently exhibits smaller deviations is considered to be more reliable, and the trust assigned to this source is updated accordingly. The concept of probability dominance formalizes this comparison.
Definition 4.4 (Probability dominance, Wrather and Yu (1982))
@@
Based on this definition, we argue that $Y\beta Z$ is a necessary condition for fully trusting one source over another in a pairwise comparison. This implies that a decision maker’s trust in source $1$ with deviations $Y$ will converge to $1$ after a finite number of events. The detailed proof of the following theorem can be found in Appendix A.4.
@@
Furthermore, one can generalize the above result to cases with multiple data sources, i.e., $H\geq 3$, stated in Theorem 4.6 below. The detailed proof is provided in Appendix A.5.
@@
Theorem 4.7 (Probability dominance and FSD, Wrather and Yu (1982))
@@
We present the proof of Theorem 4.7 in Appendix A.6. We will apply Theorem 4.7 to setup computational experiments related to probability dominance and trust update based on error later in Section 5.2.
@@ 4.4.2 Dominant trust under exponential error trust update
In this section, we establish the conditions for dominant trust when using the exponential error trust update Algorithm 2. Unlike the min-max error trust update Algorithm 1, this approach does not require probability dominance. Instead, it relies on the consistency of smaller expected prediction errors over time. The detailed proof of the following theorem can be found in Appendix A.7.
@@
for all events $i=1,2,\ldots$. Also, assume that the errors $\|\Delta\xi^{(i)}_{h}\|$ are independent over time and bounded within the interval $[0,L_{\text{max}}]$. Then, using the exponential error trust update Algorithm 2, the trust $t^{(i)}_{h_{Y}}$ will converge to $1$ as $i\rightarrow\infty$.
@@ 5 Computational Results
All linear programming reformulations of the MR-DRO model are optimized by Gurobi 9.5.2. The algorithm for trust update is implemented in Python 3.9.12. All numerical tests are conducted on a PC with 16 GB RAM and an Apple M1 Pro chip.
@@ 5.1.1 MR-DRO for a resource allocation problem
Consider a resource allocation problem in $K_{\text{r}}$ regions. Let $\bm{c}^{u}=[c^{u}_{1},\ldots,c^{u}_{K_{\text{r}}}]^{\mathsf{T}}$ be the unit penalty cost of unmet demand and $\bm{c}^{o}=[{c}^{o}_{1},\ldots,{c}^{o}_{K_{\text{r}}}]^{\mathsf{T}}$ be that of over-served demand in each region. We have an overall resource budget $B>0$. The demands in all regions are uncertain and are denoted by a random vector $\bm{d}=[d_{1},\ldots,d_{K_{\text{r}}}]^{\mathsf{T}}$, $\bm{d}\in\mathbb{R}^{K_{\text{r}}}_{+}$. We specify the decision vector $\bm{x}=[x_{1},\ldots,x_{K_{\text{r}}}]^{\mathsf{T}}$ with a feasible region $\mathbb{X}=\{\bm{x}\in\mathbb{R}^{K_{\text{r}}}_{+}:\sum_{k=1}^{K_{\text{r}}}x_{k}\leq B\}$, where $x_{k}$ indicates the amount of resource allocated to region $k$, for each $k\in[K_{\text{r}}]$. Following the generic steps of MR-DRO in Section 2.2, we solve the following problem:
@@
which minimizes the expected losses caused by undesirable allocation. Here, for a vector $\bm{u}\in\mathbb{R}^{M}$, $(\bm{u})^{+}=[\max\{u_{1},0\},\ldots,\max\{u_{M},0\}]^{\mathsf{T}}$. By expressing the loss function as the additively separable form, we obtain:
@@ 5.1.2 Baseline settings for the resource allocation problem
We consider a baseline case with $K_{\text{r}}=4$, $H=3$, and $I=200$. We set $\epsilon=0.01$ as the radius of the ambiguity set $\mathcal{P}$. The unit penalty costs for unmet and over-served demand are $\bm{c}^{u}=(5000,5000,5000)^{\mathsf{T}}$ and $\bm{c}^{o}=(1000,1000,1000)^{\mathsf{T}}$, respectively. The resource budget is $B=200$, which is sufficient for all events. The true demand $\bm{d}^{(i)}_{\textrm{true}}$ for event $i$ is a vector with each element sampled from a uniform distribution $U(10,20)$. The prediction at region $k$ from source $h$ for event $i$ is sampled from a truncated normal distribution with mean $(d^{(i)}_{\textrm{true},k}+\mu_{h,k})$ and variance $\sigma^{2}_{h,k}$ within interval $[0,30]$. Specifically, we set $\bm{\mu}=(\bm{\mu}_{1},\bm{\mu}_{2},\bm{\mu}_{3})^{\mathsf{T}}=((0,0,0,0)^{\mathsf{T}},(0,5,0,5)^{\mathsf{T}},(0,-5,5,2)^{\mathsf{T}})^{\mathsf{T}}$ and $\bm{\sigma}=(\bm{\sigma}_{1},\bm{\sigma}_{2},\bm{\sigma}_{3})^{\mathsf{T}}=((1,1,1,5)^{\mathsf{T}},(2,1,5,5)^{\mathsf{T}},(5,1,1,2)^{\mathsf{T}})^{\mathsf{T}}$. The relationship between the prediction error distributions of each source and the true realization value is demonstrated in Figure 2. The step size for min-max error trust update is $\Delta t=0.01$. The update rate is $\eta=0.5$ for exponential error trust update as well as variable-share error trust update, and the share portion is $\beta=0.01$ in variable-share trust update. We conduct $30$ trials with distinct random seeds for each trust update algorithm to observe common patterns during the trust update process.
@@ 5.1.3 Baseline results for the resource allocation problem
We also perform an out-of-sample performance analysis with the $\bm{x}^{*}(\bm{t}^{(I)})$ solved by MR-DRO models and the solutions $\bm{x}^{*}_{h}$ solved by DRO models with information from source $h$, $h\in[H]$. The out-of-sample test is conducted with an empirical distribution constructed by $|\Psi|$ events with each event $\psi\in\Psi$ having an equal probability $p^{\psi}=1/|\Psi|$. The out-of-sample performance is evaluated as:
@@ 5.1.4 Sensitivity analysis for the resource allocation problem
One of the advantages of the method to construct the ambiguity set in this study is that it does not require that the prediction errors from different information sources follow the same type of probability distribution. To test this, we modify the baseline setting such that the prediction values provided by Source $2$ and Source $3$ are sampled from Lognormal distributions instead of truncated Normal distributions (see Figure 5). The results, reported in Table 4, show that varying the distribution types does not influence the generally good performance of the MR-DRO approach.
@@
Another advantage of trust dynamics is its ability to adapt when the error distribution of each source is nonstationary, changing throughout the trust update process. In this experiment, each source’s error distribution shifts at predefined time points, reflecting real-world scenarios where the accuracy of a reference information source can fluctuate over time. We set the original mean and standard deviation of the error distribution of each source as $\bm{\mu}=(\bm{\mu}_{1},\bm{\mu}_{2},\bm{\mu}_{3})^{\mathsf{T}}=((0,0,0,0)^{\mathsf{T}},(0,5,0,0)^{\mathsf{T}},(0,-5,5,0)^{\mathsf{T}})^{\mathsf{T}}$ and $\bm{\sigma}=(\bm{\sigma}_{1},\bm{\sigma}_{2},\bm{\sigma}_{3})^{\mathsf{T}}=((1,1,5,1)^{\mathsf{T}},(2,1,2,2)^{\mathsf{T}},(5,1,1,5)^{\mathsf{T}})^{\mathsf{T}}$. The predefined time points are $(100,100,100,50)$ in each region. After the corresponding time points, $\bm{\mu}$ and $\bm{\sigma}$ change to $\bm{\mu}^{{}^{\prime}}=(\bm{\mu}_{1}^{{}^{\prime}},\bm{\mu}_{2}^{{}^{\prime}},\bm{\mu}_{3}^{{}^{\prime}})^{\mathsf{T}}=((5,0,0,5)^{\mathsf{T}},(0,5,0,5)^{\mathsf{T}},(0,-5,5,2)^{\mathsf{T}})^{\mathsf{T}}$ and $\bm{\sigma}^{{}^{\prime}}=(\bm{\sigma}_{1}^{{}^{\prime}},\bm{\sigma}_{2}^{{}^{\prime}},\bm{\sigma}_{3}^{{}^{\prime}})^{\mathsf{T}}=((1,1,5,1)^{\mathsf{T}},(2,1,2,2)^{\mathsf{T}},(5,1,1,5)^{\mathsf{T}})^{\mathsf{T}}$. For example, in Region $1$, the error distribution of source $1$ shifts from a Normal distribution $\mathcal{N}(0,1)$ to $\mathcal{N}(5,1)$ after $100$ events have occurred.
@@ 5.2 Probability dominance and dominant trust
Focusing on cases with $H=2$ and the min-max error trust update approach, we develop results regarding probability dominance and certain trust update patterns. We utilize experiments to illustrate the trust update pattern given the probability dominance relationship between two sources. We set $\bm{\mu}=(\bm{\mu}_{1},\bm{\mu}_{2})^{\mathsf{T}}=((0,0,0,2)^{\mathsf{T}},(0,5,2,-2)^{\mathsf{T}})^{\mathsf{T}}$ and $\bm{\sigma}=(\bm{\sigma}_{1},\bm{\sigma}_{2})^{\mathsf{T}}=((1,5,5,2)^{\mathsf{T}},(5,5,2,2)^{\mathsf{T}})^{\mathsf{T}}$. We observe the trust update process for $I=300$ events. Judged from the empirical cumulative distribution function and based on Theorem 4.5, the deviation from source $1$ dominates the deviation from source $2$ with probability $\beta=0.5$; while in region $3$, the deviation from source $2$ has probability dominance with $\beta=0.5$ over the deviation from source $1$. In region $4$, there is no clear probability dominance between these two sources.
@@ 5.3.1 MR-DRO for a portfolio optimization problem
Consider a capital market with $K_{\text{a}}$ assets, where the returns of these assets are characterized by the random vector $\bm{r}=[r_{1},\ldots,r_{K_{\text{a}}}]^{\mathsf{T}}$, $\bm{r}\in\mathbb{R}^{K_{\text{a}}}$. Short-selling forbidden, a portfolio can be encoded by a vector $\bm{x}=[x_{1},\ldots,x_{K_{\text{a}}}]^{\mathsf{T}}$ of percentage weights of each asset. For each asset $k\in[K_{\text{a}}]$, a proportion $x_{k}$ of the total capital is invested, leading to a portfolio return of $\langle\bm{x},\bm{r}\rangle$. In the remainder we aim to solve the the following MR-DRO model:
@@
which minimizes a weighted sum of the expectation and the conditional value at-risk (CVaR) of the portfolio loss $-\langle\bm{x},\bm{r}\rangle$, where $\alpha\in(0,1]$ is referred to as the confidence level of the CVaR, and $\rho\in\mathbb{R}_{+}$ quantifies the investor’s risk-aversion. The CVaR at the $\alpha$ level essentially captures the average of the worst $\alpha\times 100\%$ portfolio losses under distribution $\mathbb{Q}$. By expressing the CVaR with its formal definition, we arrive at
@@
where $J=2$, $\bm{a}_{1}=-\bm{x}$, $a_{2}=(-1-\rho/\alpha)\bm{x}$, $b_{1}=\rho\tau$, $b_{2}=\rho\tau(1-1/\alpha)$. In a scenario where an investor does not have access to the true distribution $\mathbb{P}$ but possesses the empirical distribution ^$\hat{\mathbb{P}}_{HI}$ constructed by non-parametric data fusion based on trust, we can apply Theorem 3.2 to determine an optimal portfolio $\bm{x}$ and formulate the problem as (18):
$\inf_{\bm{x},\lambda,s_{hi},\bm{\gamma}_{hij}}\quad$ $\lambda\epsilon+\sum^{H}_{h=1}\sum^{K}_{k=1}\sum^{I}_{i=1}\frac{t^{(I)}_{h,k}}{I}\cdot s_{hi}$ (18a)

### v2

@@ 4.2 Exponential error trust update
where $\eta>0$ is the update rate, and $\|\cdot\|$ is a norm on $\mathbb{R}^{M}$. After computing the updated trust values, they are normalized to ensure that the sum of all trust values remains constant, preserving the interpretability of trust as weights in $\hat{\mathbb{P}}_{HI}$.
@@ 4.3 Variable-share error trust update
The variable-share error trust update method extends the exponential error trust update by introducing a redistribution mechanism that dynamically adjusts trust values based on observed prediction errors and redistributes a fraction of trust among the sources \@BBOPcitep\@BAP\@BBN(Herbster and Warmuth 1998)\@BBCP. This approach balances penalizing poorly performing sources and rewarding better performing ones by incorporating a flexible share allocation component.
@@
$W^{(i)}_{\text{trust}}:=\sum_{h=1}^{H}(1-(1-\beta)^{\|\Delta\bm{\xi}^{(i)}_{h}\|})(t^{\prime})^{(i)}_{h}$,
@@ 4.4 Analyzing trust dynamics
We are particularly interested in analyzing the dynamics of trust assigned to different sources under specific trust update algorithms and understanding whether and when dominant trust exists. We focus on two methods: the min-max error trust update in Section 4.1 and the exponential error trust update in Section 4.2. We exclude the variable-share error update method in Section 4.3 from the analysis because its primary objective is to rapidly adjust trust levels when relative accuracy relationships between sources change. Consequently, the concept of dominant trust is less relevant to the variable-share error update method.
@@ 4.4.1 Dominant trust under min-max error trust update
We consider the concept of probability dominance \@BBOPcitep\@BAP\@BBN(Wrather and Yu 1982)\@BBCP to rationalize the assignment of dominant trust (if it exists) to one source over others. Let $\Omega$ denote a collection of random variables corresponding to the deviations of the predicted values from the true realizations, specifically the norm of the prediction errors. For simplicity, we first examine the case where there are only two data sources, $h_{Y}$ and $h_{Z}$, that is, $H=2$. Let $Y=\|\Delta\bm{\xi}^{(i)}_{h_{Y}}\|$ represent the norm of the prediction errors of source $h_{Y}$, and $Z=\|\Delta\bm{\xi}^{(i)}_{h_{Z}}\|$ for the source $h_{Z}$, over multiple events $i=1,2,\ldots$. The cumulative distribution functions of $Y$ and $Z$ are well defined and are denoted by $F_{Y}(a)$ and $F_{Z}(a)$, respectively. In this framework, a source that consistently exhibits smaller deviations is considered to be more reliable, and the trust assigned to this source is updated accordingly. The concept of probability dominance formalizes this comparison.
Definition 4.4 (Probability dominance, \@BBOPcitet\@BAP\@BBNWrather and Yu (1982)\@BBCP)
@@
Based on this definition, we argue that $Y\beta Z$ is a necessary condition for fully trusting one source over another in a pairwise comparison. This implies that a decision maker’s trust in source $1$ with deviations $Y$ will converge to $1$ after a finite number of events. The detailed proof of the following theorem can be found in Supplement Material Section A.4.
@@
Furthermore, one can generalize the above result to cases with multiple data sources, i.e., $H\geq 3$, stated in Theorem 4.6 below. The detailed proof is provided in Supplement Material Section A.5.
@@
Theorem 4.7 (Probability dominance and FSD, \@BBOPcitet\@BAP\@BBNWrather and Yu (1982)\@BBCP)
@@
We present the proof of Theorem 4.7 in Supplement Material Section A.6.
@@ 4.4.2 Dominant trust under exponential error trust update
In this section, we establish the conditions for dominant trust when using the exponential error trust update Algorithm 2. Unlike the min-max error trust update Algorithm 1, this approach does not require probability dominance. Instead, it relies on the consistency of smaller expected prediction errors over time. The detailed proof of the following theorem can be found in Supplement Material Section A.7.
@@
for all events $i=1,2,\ldots$. Also, assume that the errors $\|\Delta\xi^{(i)}_{h}\|$ are independent over time and bounded within the interval $0,L_{\text{max}}]$. Then, using the exponential error trust update Algorithm 2, the trust $t^{(i)}_{h_{Y}}$ will converge to $1$ as $i\rightarrow\infty$.
@@ 5 Computational Results
All linear programming models (i.e., reformulations of the MR-DRO model) are optimized by Gurobi 9.5.2. The algorithm for trust update is implemented in Python 3.9.12. All numerical tests are conducted on a PC with 16 GB RAM and an Apple M1 Pro chip.
@@ 5.1.1 MR-DRO for a resource allocation problem
We optimize the allocation of resources in $K_{\text{r}}$ regions, constrained by an overall budget $B>0$. The demands $\bm{d}=[d_{1},\ldots,d_{K_{\text{r}}}]^{\mathsf{T}}$, $\bm{d}\in\mathbb{R}^{K_{\text{r}}}_{+}$ in all regions are uncertain and they can be estimated using predictions from multiple $h=1,\ldots,H$ sources. Let $\bm{c}^{u}=[c^{u}_{1},\ldots,c^{u}_{K_{\text{r}}}]^{\mathsf{T}}$ be the unit penalty cost of unmet demand and $\bm{c}^{o}=[{c}^{o}_{1},\ldots,{c}^{o}_{K_{\text{r}}}]^{\mathsf{T}}$ be that of over-served demand in each region. We specify the decision vector $\bm{x}=[x_{1},\ldots,x_{K_{\text{r}}}]^{\mathsf{T}}$ with a feasible region $\mathbb{X}=\{\bm{x}\in\mathbb{R}^{K_{\text{r}}}_{+}:\sum_{k=1}^{K_{\text{r}}}x_{k}\leq B\}$, where $x_{k}$ indicates the amount of resource allocated to region $k$, for each $k\in[K_{\text{r}}]$. For any vector $\bm{u}\in\mathbb{R}^{M}$, we define the function $(\bm{u})^{+}:=[\max\{u_{1},0\},\ldots,\max\{u_{M},0\}]^{\mathsf{T}}$. Following the structure of MR-DRO given in Section 2.2, we formulate:
@@
which minimizes the expected losses caused by undesirable allocation. By expressing the loss function as the additively separable form, we obtain:
@@ 5.1.2 Baseline settings for the resource allocation problem
We consider a baseline case with $K_{\text{r}}=4$, $H=3$, and $I=200$. We set $\epsilon=0.01$ as the radius of the ambiguity set $\mathcal{P}$. The unit penalty costs for unmet and over-served demand are $\bm{c}^{u}=(\mathdollar 5000,\mathdollar 5000,\mathdollar 5000,\mathdollar 5000)^{\mathsf{T}}$ and $\bm{c}^{o}=(\mathdollar 1000,\mathdollar 1000,\mathdollar 1000,\mathdollar 1000)^{\mathsf{T}}$, respectively. The resource budget is $B=200$. The true demand $\bm{d}^{(i)}_{\textrm{true}}$ for event $i$ is a vector with each element sampled from a uniform distribution $U(10,20)$. The prediction at region $k$ from source $h$ for event $i$ is sampled from a truncated normal distribution with mean $(d^{(i)}_{\textrm{true},k}+\mu_{h,k})$ and variance $\sigma^{2}_{h,k}$ within interval $[0,30]$. Specifically, we set $\bm{\mu}=(\bm{\mu}_{1},\bm{\mu}_{2},\bm{\mu}_{3})^{\mathsf{T}}=((0,0,0,0)^{\mathsf{T}},(0,5,0,5)^{\mathsf{T}},(0,-5,5,2)^{\mathsf{T}})^{\mathsf{T}}$ and $\bm{\sigma}=(\bm{\sigma}_{1},\bm{\sigma}_{2},\bm{\sigma}_{3})^{\mathsf{T}}=((1,1,5,5)^{\mathsf{T}},(2,1,1,5)^{\mathsf{T}},(5,1,1,2)^{\mathsf{T}})^{\mathsf{T}}$. The relationship between the prediction error distributions of each source and the true realization value is demonstrated in Figure 2. The step size for min-max error trust update is $\Delta t=0.01$. The update rate is $\eta=0.5$ for both exponential error trust update and variable-share error trust update methods, and the share portion is $\beta=0.01$ in the variable-share trust update method. We conduct $30$ trials with distinct random seeds for each trust update algorithm.
@@ 5.1.3 Baseline results for the resource allocation problem
We also perform an out-of-sample performance analysis with the $\bm{x}^{*}(\bm{t}^{(I)})$ solved by MR-DRO models and the solutions $\bm{x}^{*}_{h}$ solved by DRO models with information from source $h$, $\forall h\in[H]$. The out-of-sample test is conducted with an empirical distribution constructed by $|\Psi|$ events with each event $\psi\in\Psi$ having an equal probability $p^{\psi}=1/|\Psi|$. The out-of-sample performance is evaluated as:
@@ 5.1.4 Sensitivity analysis for the resource allocation problem
One of the advantages of our approach that constructs the ambiguity set in this study is that it does not require that the prediction errors from different information sources follow the same type of probability distribution. To test this, we modify the baseline setting such that the prediction values provided by Source $2$ and Source $3$ are sampled from Lognormal distributions instead of truncated Normal distributions (see Figure 5). The results, reported in Table 4, show that varying the distribution types does not influence the performance of MR-DRO, and it still outperforms other benchmark models that rely on a single information source.
@@
Another advantage of trust dynamics is its ability to adapt when the error distribution of each source is nonstationary, changing throughout the trust update process. In this experiment, each source’s error distribution shifts at predefined time points, reflecting real-world scenarios where the accuracy of a reference information source can fluctuate over time. We set the original mean and standard deviation of the error distribution of each source as $\bm{\mu}=(\bm{\mu}_{1},\bm{\mu}_{2},\bm{\mu}_{3})^{\mathsf{T}}=((0,0,0,0)^{\mathsf{T}},(0,5,0,5)^{\mathsf{T}},(0,-5,5,2)^{\mathsf{T}})^{\mathsf{T}}$ and $\bm{\sigma}=(\bm{\sigma}_{1},\bm{\sigma}_{2},\bm{\sigma}_{3})^{\mathsf{T}}=((1,1,5,5)^{\mathsf{T}},(2,1,1,5)^{\mathsf{T}},(5,1,1,2)^{\mathsf{T}})^{\mathsf{T}}$. The predefined time points are $(100,100,100,50)$ in each region. After the corresponding time points, $\bm{\mu}$ and $\bm{\sigma}$ change to $\bm{\mu}^{{}^{\prime}}=(\bm{\mu}_{1}^{{}^{\prime}},\bm{\mu}_{2}^{{}^{\prime}},\bm{\mu}_{3}^{{}^{\prime}})^{\mathsf{T}}=((5,0,0,5)^{\mathsf{T}},(0,0,0,0)^{\mathsf{T}},(0,-5,0,0)^{\mathsf{T}})^{\mathsf{T}}$ and $\bm{\sigma}^{{}^{\prime}}=(\bm{\sigma}_{1}^{{}^{\prime}},\bm{\sigma}_{2}^{{}^{\prime}},\bm{\sigma}_{3}^{{}^{\prime}})^{\mathsf{T}}=((1,1,5,1)^{\mathsf{T}},(2,1,2,2)^{\mathsf{T}},(5,1,1,5)^{\mathsf{T}})^{\mathsf{T}}$. For example, in Region $1$, the error distribution of source $1$ shifts from a Normal distribution $\mathcal{N}(0,1)$ to $\mathcal{N}(5,1)$ after $100$ events have occurred.
@@ 5.2 Probability dominance and dominant trust
In this section, we numerically validate the result of probability dominance in Theorem 4.5 and illustrate the related trust update patterns, by focusing on special cases with $H=2$ information sources. For $K_{r}=4$ regions, we set $\bm{\mu}=(\bm{\mu}_{1},\bm{\mu}_{2})^{\mathsf{T}}=((0,0,0,2)^{\mathsf{T}},(0,5,2,-2)^{\mathsf{T}})^{\mathsf{T}}$ and $\bm{\sigma}=(\bm{\sigma}_{1},\bm{\sigma}_{2})^{\mathsf{T}}=((1,5,5,2)^{\mathsf{T}},(5,5,2,2)^{\mathsf{T}})^{\mathsf{T}}$. We observe the trust update process for $I=300$ events. Judged from the empirical cumulative distribution function and based on Theorem 4.5, the deviation from source $1$ dominates the deviation from source $2$ with probability $\beta=0.5$; while in region $3$, the deviation from source $2$ has probability dominance with $\beta=0.5$ over the deviation from source $1$. In region $4$, there is no clear probability dominance between these two sources.
@@ 5.3 Portfolio optimization with uncertain return
We provide another set of numerical studies based on portfolio optimization instances, where the uncertainty appears in the objective function.
@@ 5.3.1 MR-DRO for a portfolio optimization problem
Consider a capital market with $K_{\text{a}}$ assets, where the returns of these assets are characterized by the random vector $\bm{r}=[r_{1},\ldots,r_{K_{\text{a}}}]^{\mathsf{T}}$, $\bm{r}\in\mathbb{R}^{K_{\text{a}}}$. Short-selling forbidden, a portfolio can be encoded by a vector $\bm{x}=[x_{1},\ldots,x_{K_{\text{a}}}]^{\mathsf{T}}$ of percentage weights of each asset. For each asset $k\in[K_{\text{a}}]$, a proportion $x_{k}$ of the total capital is invested, leading to a portfolio return of $\langle\bm{x},\bm{r}\rangle$. Accordingly, we solve the the following MR-DRO model:
@@
which minimizes a weighted sum of the expectation and the conditional value at-risk (CVaR) of the portfolio loss $\langle\bm{x},\bm{r}\rangle$, where $\alpha\in(0,1]$ is referred to as the confidence level of the CVaR, and $\rho\in\mathbb{R}_{+}$ quantifies the investor’s risk-aversion. The CVaR at the $\alpha$ level essentially captures the average of the worst $\alpha\times 100\%$ portfolio losses under distribution $\mathbb{Q}$. By expressing the CVaR with its formal definition, we arrive at
@@
where $J=2$, $\bm{a}_{1}=-\bm{x}$, $a_{2}=(-1-\rho/\alpha)\bm{x}$, $b_{1}=\rho\tau$, $b_{2}=\rho\tau(1-1/\alpha)$. In a scenario where an investor does not have access to the true distribution $\mathbb{P}$ but possesses the empirical distribution $\hat{\mathbb{P}}_{HI}$ constructed by nonparametric data fusion based on trust, we can apply Theorem 3.2 to determine an optimal portfolio $\bm{x}$ and formulate the problem as (18):
$\inf_{\bm{x},\lambda,s_{hi},\bm{\gamma}_{hij}}\quad$ $\lambda\epsilon+\sum_{h=1}^{H}\sum_{i=1}^{I}\frac{t^{(I)}_{h}}{I}s_{hi}$ (18a)